# PRISM: A Framework Harnessing Unsupervised Visual Representations and Textual Prompts for Explainable MACE Survival Prediction from Cardiac Cine MRI

### v1

@@ Abstract
Accurate prediction of major adverse cardiac events (MACE) remains a central challenge in cardiovascular prognosis. We present PRISM (Prompt-guided Representation Integration for Survival Modeling), a self-supervised framework that integrates visual representations from non-contrast cardiac cine magnetic resonance imaging with structured electronic health records (EHRs) for survival analysis. PRISM extracts temporally synchronized imaging features through motion-aware multi-view distillation and modulates them using medically informed textual prompts to enable fine-grained risk prediction. Across four independent clinical cohorts, PRISM consistently surpasses classical survival prediction models and state-of-the-art (SOTA) deep learning baselines under internal and external validation. Further clinical findings demonstrate that the combined imaging and EHR representations derived from PRISM provide valuable insights into cardiac risk across diverse cohorts. Three distinct imaging signatures associated with elevated MACE risk are uncovered, including lateral wall dyssynchrony, inferior wall hypersensitivity, and anterior elevated focus during diastole. Prompt-guided attribution further identifies hypertension, diabetes, and smoking as dominant contributors among clinical and physiological EHR factors. The source code is available at https://github.com/Hoyant-Su/PRISM.
Keywords: Major adverse cardiac events, Non-contrast survival prediction, Motion-aware multi-view distillation, Prompt-guided medical analysis
@@ 1 Introduction
Major adverse cardiovascular events (MACE), encompassing outcomes such as cardiovascular death, myocardial infarction, and hospitalization for heart failure, remain a significant cause of morbidity and mortality globally [1, 2, 3, 4]. As a composite endpoint, accurate prediction of MACE is essential for informing clinical decisions, optimizing patient outcomes, and capturing clinically significant deterioration, which correlates with long-term prognosis. This makes MACE prediction central to cardiovascular risk stratification [5, 6]. Consequently, many predictive approaches, such as AI-driven methods [7, 8, 9, 10, 11, 12, 13, 4, 14, 15], have become increasingly critical in MACE prediction in recent years.
Traditional approaches [16, 17, 18] often purely rely on manually curated features derived from electronic health records (EHRs), which, despite their accessibility, offer refined but limited physiological resolution and may fail to reflect the dynamic nature of cardiac function. Alternatively, methods based on histopathological images [19, 20, 17, 21, 22], while providing high-resolution cellular information, are inherently static and dimensionally restricted, making them less suitable for capturing organ-level or temporal characteristics relevant to adverse cardiovascular outcomes. These limitations hinder the extraction of comprehensive prognostic representations. In contrast, cardiac cine magnetic resonance imaging (MRI) provides organ-scale, high-resolution, and temporally dynamic information [23, 24], enabling the discovery of more expressive biomarkers associated with MACE [25, 26, 27]. In this context, as Fig. 1 shows, we exploit the complementary strengths of cine MRI and structured EHR data to construct a unified multimodal dataset tailored for integrative prognostic modeling.
@@
(a) PRISM’s three-stage survival analysis framework.
(b) Analysis pipeline based on PRISM.
@@
Figure 2: Details of the proposed framework and pipeline. (a): PRISM’s three-stage survival analysis framework. Multi-view cine MRI images passed through a teacher-student distillation network to obtain representation tokens enriched with spatiotemporal features, aligned hierarchically with EHR features under the influence of medically informed prompt aggregation. The refined image features are ultimately aggregated with EHR features in Stage III to yield survival analysis results. (b): Medical insights discovery pipeline based on PRISM. Deriving the association between spatiotemporal patterns of ventricular myocardium and MACE risk from the learned heatmap distributions, and exploring EHR features aiding in MACE assessment based on the BiPromptSurv strategy.
@@ 2.1 Generalizability of PRISM across internal–external validation settings in multi-center cohorts
To rigorously evaluate the robustness and generalizability of our proposed framework, we conducted a comprehensive internal–external cross-validation (IECV) [33] analysis across four independent cohorts. PRISM was comparatively assessed against classical statistical models, neural network–based approaches, and contemporary state-of-the-art (SOTA) survival methodologies using the concordance index (C-index) and time-dependent area under the curve (AUC) as primary performance metrics. The comparator models encompassed the Cox proportional hazards model (CoxPH) [18], established as the clinical gold standard for survival analysis, as well as DeepSurv [17] and Deep Survival Machines (DSM) [16], designed to capture non-linear risk functions from tabular EHR data. Represented SOTA models in 3D medical imaging–based survival analysis, including Sparse BagNet and SurvRNC, were utilized to benchmark the performance of PRISM. Furthermore, to evaluate cross-paradigm efficacy within unsupervised representation learning, PRISM was benchmarked against two prominent self-supervised frameworks, namely PCRL [30] and 3DINO [28], which pretrained on large-scale, multimodal medical imaging datasets.
As illustrated in Fig. 3a and Fig. 3b, PRISM consistently demonstrated superior performance in survival analysis across both in-domain and out-of-domain validations, with its boxplot distributions exhibiting higher medians and narrower interquartile ranges compared to baseline models. This pattern was especially evident in the GLCCM and AZCCM cohorts, where alternative methods showed either lower central tendencies or pronounced variability across seeds. The stability of PRISM’s performance distributions indicates its robustness, thereby enhancing the trustworthiness of prognostic assessments and supporting more precise decision-making.
Compared to models rooted in clinical tabular modeling paradigms such as CoxPH, DeepSurv, and DSM, PRISM demonstrated marked gains particularly in scenarios characterized by multimodal feature heterogeneity, highlighting that image-driven alignment of latent EHR prognostic factors underpins its enhanced effectiveness. While traditional models generally exhibited sharp performance decay in out-domain settings in the RJCCM and TJCCM cohorts, PRISM retained stable performance under distribution shift.
Furthermore, when benchmarked against advanced SSL frameworks pretrained on large-scale medical imaging corpora, including PCRL and 3DINO, PRISM consistently exhibited higher discriminative capacity across both low- and high-risk strata.
(a) Boxplots of C-index and AUC measured under in-domain conditions.
(b) Boxplots of C-index and AUC measured under out-domain conditions.
(c) Regression of predicted risk against ground-truth survival time.
Figure 3: Model survival analysis performance under the IECV setting. (a), (b): Boxplots of C-index (amber) and AUC (teal) across five data-splitting strategies under internal and external validation settings, respectively. (c): Regression analysis in the interval-external cross-validation setting, with model-predicted inverse risk (normalized) on the horizontal axis versus ground-truth survival time (normalized) on the vertical axis. Top panels represent internal validation and bottom panels show external validation in correlated cohorts. Cases with MACE=0 are denoted in black, while those with MACE=1 are indicated in blue.
Fig. 3c demonstrates the linear relationship between log-transformed inverse risk scores from our prognostic model and observed survival times in right-censored specimens. This further confirms the discriminative efficacy of computed risk in characterizing survival distributions. Specimens with MACE=1 consistently occupy high-risk positions across test sets in IECV setting. Critically, validation under external-center training conditions manifested only a minor slope reduction in average relative to internal validation. Specifically, the confidence band in the external GLCCM cohort exhibits progressive convergence with extended survival times, yielding statistically narrower 95% confidence intervals versus the green internal band. This demonstrates time-invariant precision in survival estimation and temporal stability within the IECV experimental setting.
Corresponding to Fig. 3a and Fig. 3b, Table 1 and Table 2 present the detailed comparison results, including the ablation experiments. PRISM without Stage II represents an ablation version of our proposed model in which the visual features from Stage I are directly fed into survival analysis. The best and second-best performances for each cohort are highlighted in bold and underlined, respectively. It can be observed that PRISM demonstrates the highest average performance across the four cohorts, ranking first or second in most, and remaining highly competitive throughout. To further evaluate the significance of these results across random seeds, we combined the p-values obtained from each seed using Fisher’s method [34]. The test statistic was computed as $X^{2}=-2\sum_{i=1}^{k}\ln(p_{i})$, which follows a $\chi^{2}$ distribution with $2k$ degrees of freedom.
@@
CoxPH 0.698±0.071 0.760±0.020 0.568±0.040 0.619±0.091
DSM 0.572±0.079 0.656±0.048 0.536±0.076 0.532±0.037
DeepSurv 0.553±0.095 0.616±0.092 0.545±0.104 0.593±0.087
SurvRNC 0.641±0.091 0.716±0.063 0.583±0.109 0.579±0.043
Sparse BagNet 0.536±0.050 0.547±0.110 0.532±0.146 0.495±0.102
3DINO 0.573±0.047 0.588±0.049 0.614±0.055 0.634±0.115
PCRL 0.701±0.052 0.749±0.044 0.526±0.089 0.649±0.098
PRISM (Ours, w/o Stage II) 0.600±0.031 0.671±0.049 0.615±0.072 0.730±0.079
PRISM (Ours) 0.690±0.055 0.772±0.035 0.617±0.033 0.775±0.037
@@
CoxPH 0.708±0.055 0.671±0.076 0.599±0.039 0.499±0.081
DSM 0.575±0.038 0.621±0.054 0.463±0.101 0.472±0.111
DeepSurv 0.559±0.125 0.579±0.171 0.553±0.139 0.565±0.160
SurvRNC 0.686±0.079 0.717±0.044 0.582±0.038 0.513±0.093
Sparse BagNet 0.563±0.077 0.459±0.040 0.492±0.077 0.559±0.103
3DINO 0.587±0.030 0.566±0.050 0.581±0.080 0.581±0.120
PCRL 0.714±0.049 0.609±0.046 0.605±0.053 0.573±0.021
PRISM (Ours, w/o Stage II) 0.617±0.061 0.679±0.047 0.665±0.088 0.608±0.115
PRISM (Ours) 0.710±0.010 0.712±0.031 0.609±0.077 0.701±0.075
@@
CoxPH 0.693±0.083 0.709±0.042 0.574±0.049 0.601±0.074
DSM 0.562±0.060 0.626±0.058 0.520±0.065 0.521±0.038
DeepSurv 0.543±0.105 0.583±0.070 0.526±0.112 0.574±0.095
SurvRNC 0.645±0.075 0.682±0.077 0.583±0.121 0.552±0.061
Sparse BagNet 0.507±0.054 0.562±0.084 0.542±0.130 0.509±0.140
3DINO 0.586±0.049 0.558±0.039 0.620±0.052 0.600±0.119
PCRL 0.699±0.075 0.699±0.051 0.500±0.087 0.657±0.092
PRISM (Ours, w/o Stage II) 0.606±0.052 0.667±0.055 0.592±0.068 0.744±0.081
PRISM (Ours) 0.689±0.036 0.714±0.016 0.593±0.030 0.760±0.066
@@
CoxPH 0.669±0.072 0.676±0.067 0.609±0.033 0.474±0.076
DSM 0.569±0.026 0.606±0.055 0.479±0.079 0.463±0.114
DeepSurv 0.552±0.108 0.599±0.130 0.551±0.118 0.559±0.132
SurvRNC 0.683±0.073 0.678±0.050 0.571±0.059 0.481±0.095
Sparse BagNet 0.552±0.087 0.426±0.016 0.505±0.070 0.559±0.135
3DINO 0.595±0.012 0.547±0.043 0.574±0.075 0.600±0.145
PCRL 0.713±0.048 0.582±0.042 0.594±0.051 0.544±0.035
PRISM (Ours, w/o Stage II) 0.622±0.057 0.651±0.059 0.663±0.092 0.568±0.148
PRISM (Ours) 0.692±0.019 0.676±0.028 0.589±0.091 0.657±0.130

### v2

@@ Abstract
Accurate prediction of major adverse cardiac events (MACE) remains a central challenge in cardiovascular prognosis. We present PRISM (Prompt-guided Representation Integration for Survival Modeling), a self-supervised framework that integrates visual representations from non-contrast cardiac cine magnetic resonance imaging with structured electronic health records (EHRs) for survival analysis. PRISM extracts temporally synchronized imaging features through motion-aware multi-view distillation and modulates them using medically informed textual prompts to enable fine-grained risk prediction. Across four independent clinical cohorts, PRISM consistently surpasses classical survival prediction models and state-of-the-art (SOTA) deep learning baselines under internal and external validation. Further clinical findings demonstrate that the combined imaging and EHR representations derived from PRISM provide valuable insights into cardiac risk across diverse cohorts. Three distinct imaging signatures associated with elevated MACE risk are uncovered, including lateral wall dyssynchrony, inferior wall hypersensitivity, and anterior elevated focus during diastole. Prompt-guided attribution further identifies hypertension, diabetes, and smoking as dominant contributors among clinical and physiological EHR factors.
keywords:
Major adverse cardiac events , Non-contrast survival prediction , Motion-aware multi-view distillation , Prompt-guided medical analysis
journal: IEEE Transactions on Medical Imaging
[…]
[label1]organization=Fudan University, city=Shanghai, country=China
[…]
[label2]organization=Shanghai Innovation Institute, city=Shanghai, country=China
[…]
[label3]organization=Shanghai Artificial Intelligence Laboratory, city=Shanghai, country=China
[…]
[label4]organization=Shanghai Jiao Tong University, city=Shanghai, country=China
[…]
[label5]organization=University of Science and Technology of China, city=Hefei, country=China
@@ 1 Introduction
Major adverse cardiovascular events (MACE), encompassing outcomes such as cardiovascular death, myocardial infarction, and hospitalization for heart failure, remain a significant cause of morbidity and mortality globally [47, 28, 36, 53]. As a composite endpoint, accurate prediction of MACE is essential for informing clinical decisions, optimizing patient outcomes, and capturing clinically significant deterioration, which correlates with long-term prognosis. This makes MACE prediction central to cardiovascular risk stratification [1, 35]. Consequently, many predictive approaches, such as AI-driven methods [39, 34, 38, 13, 40, 8, 44, 53, 18, 55, 27], have become increasingly critical in MACE prediction in recent years.
Traditional approaches [30, 29, 5] often purely rely on manually curated features derived from electronic health records (EHRs), which, despite their accessibility, offer refined but limited physiological resolution and may fail to reflect the dynamic nature of cardiac function. Alternatively, methods based on histopathological images [45, 46, 29, 48, 3], while providing high-resolution cellular information, are inherently static and dimensionally restricted, making them less suitable for capturing organ-level or temporal characteristics relevant to adverse cardiovascular outcomes. These limitations hinder the extraction of comprehensive prognostic representations. In contrast, cardiac cine magnetic resonance imaging (MRI) provides organ-scale, high-resolution, and temporally dynamic information [33, 20, 15, 11], enabling the discovery of more expressive biomarkers associated with MACE [26, 21, 17]. In this context, as Fig. 1 shows, we exploit the complementary strengths of cine MRI and structured EHR data to construct a unified multimodal dataset tailored for integrative prognostic modeling.
@@
Figure 2: Design of the proposed framework and pipeline. (a), PRISM’s three-stage survival analysis framework. Multi-view cine MRI images passed through a teacher-student distillation network to obtain representation tokens enriched with spatiotemporal features, aligned hierarchically with EHR features under the influence of medically informed prompt aggregation. The refined image features are ultimately aggregated with EHR features in Stage III to yield survival analysis results. (b), Medical insights discovery pipeline based on PRISM. Deriving the association between spatiotemporal patterns of ventricular myocardium and MACE risk from the learned heatmap distributions, and exploring EHR features aiding in MACE assessment based on the BiPromptSurv strategy.
2 Materials and methods
2.1 Image preprocessing: ventricular motion focus
To rigorously quantify myocardial dynamics across consecutive cine frames, we employed the Farnebäck optical flow algorithm [10] on full cardiac volumes $\mathcal{V}^{(total)}$. This approach computes dense displacement fields that provide a continuous and high-resolution representation of local tissue motion, thereby capturing the spatiotemporal behavior of ventricular wall kinematics. The resulting displacement vector field was formulated as
where $\mathbf{F}^{(t)}$ denotes the motion vector field between successive frames $\mathcal{I}_{t}$ and $\mathcal{I}_{t+1}$, and $\Psi_{\mathrm{FB}}$ represents the Farnebäck optical flow operator. This formulation allows for voxel-level characterization of inter-frame myocardial motion, which is subsequently refined by focusing on a standardized region-of-interest centered on the ventricular chambers with a standardized 96-pixel window encompassing the cardiac chambers, thereby focusing the analysis on ventricular myocardial motion while excluding signals from extracardiac structures.
2.2 Stage I: motion-aware multi-view distillation
To achieve dimensional alignment with LAX, the SAX images are processed in a slice-wise manner by the Spatial Aggregation Module as Fig. 3 shows, yielding latent representations. To capture comprehensive myocardial motion patterns, the framework implements a teacher-student architecture inspired by DINOv2 [32], where the backbone is inherited from UniFormer [25] for temporal modeling, as illustrated in Fig. 3. The teacher pathway processes long-axis views comprising 2-chamber, 3-chamber, and 4-chamber orientations, generating multi-view feature embeddings that encapsulate global myocardial dynamics. The student pathway encodes spatially compressed short-axis sequences to extract condensed motion representations. Knowledge transfer is enforced by the distillation loss,
where $\bar{p}^{(t)}$ represents the teacher’s multi-view feature embeddings, $p^{(s)}$ denotes the student’s short-axis representations, and $D_{\text{KL}}$ is the Kullback-Leibler divergence [23]. The complementary contrastive term $\mathcal{L}_{\text{contrastive}}$, implemented via InfoNCE [2], enforces intra-patient phase-specific feature consistency while separating motion characteristics across different subjects. This formulation yields spatiotemporally coherent motion tokens that capture prognostic signatures of ventricular dysfunction and facilitate downstream modeling.
Figure 3: Details of model architecture and sub-modules. The Spatial Aggregation Module aggregates latent representations obtained from layer-wise convolutions over cardiac cine MRI sequences. The Semantic Encoder encodes medical guided prompts $\mathcal{T}$ based on pretrained embeddings. The Motion-Aware Encoder incorporates the CBlock and SABlock modules from the UniFormer [25] backbone.
2.3 Stage II: EHR-attention directive
As Fig. 3 shows, to enable effective cross-modal alignment between clinical and imaging data, the model employs a routing module $\mathcal{R}$ that maps a medical prompt $\mathcal{T}\in\mathbb{V}^{L}$, where $\mathbb{V}$ denotes the clinical vocabulary and $L$ is the sequence length, into a sparse group attention vector,
$\boldsymbol{\alpha}=\mathcal{R}(\text{Embed}(\mathcal{T})),$ (3)
where each $\alpha_{i}$ is approximately 0, 0.5, or 1.
The architecture then branches into three parallel streams to establish cross-modal alignment. The Cross-Attention Fusion stream constructs query projections $\mathbf{Q}=\mathbf{P}\mathbf{W}_{Q}$ from BERT-encoded prompt embeddings [22] and linearly projects spatial-temporal visual features $\mathbf{Z}^{(s)}$ to obtain keys and values $\mathbf{K}=\mathbf{V}=\mathbf{Z}^{(s)}\mathbf{W}_{K}$. The aligned image representation is computed as
where $\oplus$ denotes concatenation and $\mathbb{E}_{l}[\cdot]$ represents average pooling over prompt tokens. For EHR feature processing, each categorical variable $e_{f}$ in group $g$ is embedded via dedicated layers $\phi_{f}(\cdot)$ and aggregated with routing weights $\alpha_{g}$,
where $S_{g}$ indexes the clinical feature subgroup. Finally, the Visual Anchor stream preserves the original spatial-temporal feature topology via
$\mathbf{Z}_{ref}=\text{LayerNorm}\left(\mathbf{W}_{r}\cdot\mathbb{E}[\mathbf{Z}^{(s)}]\right),$ (6)
ensuring structural integrity while supporting subsequent alignment objectives.
2.4 Alignment Objectives
where $B$ denotes the batch size. Then we obtain the composite objective in Equ. 9,
$\mathcal{L}_{total}=\mathcal{L}_{tri}+\beta\mathcal{L}_{pres}.$ (9)
2.5 Stage III: survival prediction
To predict patient survival, clinical indicators $\mathbf{e}\in\mathbb{R}^{d_{m}}$ and pre-trained image tokens $\mathbf{r}=\mathbb{E}(Q_{img})\in\mathbb{R}^{d_{c}}$ are concatenated into a fused representation,
$\mathbf{x}_{\mathrm{fused}}=\mathrm{concat}[\mathbf{e},\mathbf{r}]\in\mathbb{R}^{d_{m}+d_{c}},$ (10)
integrating information from both clinical and imaging domains. A Cox proportional hazards (CoxPH) head is applied, where the coefficients $\beta_{k}$ automatically weight the contribution of each fused feature. The hazard function is expressed as
with $h_{0}(t)$ representing the baseline hazard function. Training minimizes the negative log-partial likelihood,
where $\mathbf{\theta}=[\beta_{1},\ldots,\beta_{m}]^{\top}$ denotes the regression coefficients, $\delta_{i}$ indicates whether the event occurred for patient $i$, and $R(t_{i})$ corresponds to the risk set at survival time $t_{i}$. This provides a principled framework to integrate clinical and imaging data for prognostic assessment.
2.6 Evaluation metrics
In this study, we utilized C-index [12] and area under the curve (AUC) as primary evaluation metrics. The C-index assesses the discriminatory power of the model by measuring the proportion of all comparable subject pairs whose predicted risks are correctly ordered relative to their observed survival times. Formally, it is defined in Equ. 13,
where $\delta_{i}$ indicates the occurrence of the event of interest, and $r_{i}=\mathbf{\beta}^{\mathrm{T}}\mathrm{x}_{\mathrm{fused}}^{i}$ represents the risk score derived from the prognostic model. Complementarily, we incorporated AUC to quantify the model’s sensitivity and specificity in stratifying patients with respect to event status across varying decision thresholds. The time-dependent AUC at time $t$ is defined as
$\text{AUC}(t)=P(r_{i}>r_{j}|y_{i}(t)=1,y_{j}(t)=0),$ (14)
where $y_{i}(t)$ denotes the event status of subject $i$ at time $t$, and $r_{i}$ represents the predicted risk score. This metric measures the probability that a subject experiencing an event at time $t$ has a higher risk score than a subject who remains event-free at that time point.
2.7 Implementation details
The model was implemented in the PyTorch environment and trained using NVIDIA-H100 GPUs with CUDA 12.2. Specifically, Stages I and III were conducted on a single H100 GPU, whereas Stage II utilized a 4-GPU H100 setup. The GPU-to-CPU ratio was maintained at 1 to 15. Optimization was performed using the Adam optimizer [19] with an initial learning rate of $5\times 10^{-5}$ and weight decay set to $1\times 10^{-5}$. A batch size of 16 was used throughout, and models were trained for 50 epochs with a StepLR learning rate scheduler. The 4D cardiac cine MRI inputs were uniformly resized to dimensions of $24\times 24\times 96\times 96$, indicating depth, temporal frames, height, and width. A total of 41 EHR features were incorporated, with 579, 672, 406, and 313 samples included from GLCCM, AZCCM, RJCCM, and TJCCM, respectively.
Standard data partitioning schemes were employed, with a training-to-validation ratio of 7:3 for self-supervised learning during Stages I and II, and a split of 6:2:2 for training, validation, and testing sets in the survival prediction stage. In Stage II, each image sample was paired with 50 unique medical prompt variants. During survival prediction stage, grid search combined with stratified sampling was used for hyperparameter tuning, and feature selection was performed using stepwise regression and Lasso regularization within the CoxPH. The internal–external cross-validation (IECV) [41] setting is implemented by fixing the same test set within each cohort. For internal validation, training and testing are performed on the individual cohort separately. For external validation, the model is trained on the combined training sets of the remaining cohorts and tested on the same test set as used in the internal validation.
@@ 3.1 Generalizability of PRISM across internal–external validation settings in multi-center cohorts
To rigorously evaluate the robustness and generalizability of our proposed framework, we conducted a comprehensive IECV analysis across four independent cohorts. PRISM was comparatively assessed against classical statistical models, neural network–based approaches, and contemporary SOTA survival methodologies using the concordance index (C-index) and time-dependent AUC as primary performance metrics. The comparator models encompassed the CoxPH [5], established as the clinical gold standard for survival analysis, as well as DeepSurv [29] and Deep Survival Machines (DSM) [30], designed to capture non-linear risk functions from tabular EHR data. Represented SOTA models in 3D medical imaging–based survival analysis, including Sparse BagNet and SurvRNC, were utilized to benchmark the performance of PRISM. Furthermore, to evaluate cross-paradigm efficacy within unsupervised representation learning, PRISM was benchmarked against two prominent self-supervised frameworks, namely PCRL [56] and 3DINO [52], which pretrained on large-scale, multimodal medical imaging datasets.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Table 1 and Table 2 present the detailed comparison results, with the best and second-best performances for each cohort are highlighted in bold and underlined, respectively. PRISM consistently demonstrated superior performance in survival analysis across both in-domain and out-of-domain validations. Compared to models rooted in clinical tabular modeling paradigms such as CoxPH, DeepSurv, and DSM, PRISM demonstrated marked gains particularly in scenarios characterized by multimodal feature heterogeneity, highlighting that image-driven alignment of latent EHR prognostic factors underpins its enhanced effectiveness. While traditional models generally exhibited sharp performance decay in out-domain settings in the RJCCM and TJCCM cohorts, PRISM retained stable performance under distribution shift. Furthermore, when benchmarked against advanced SSL frameworks pretrained on large-scale medical imaging corpora, including PCRL and 3DINO, PRISM consistently exhibited higher discriminative capacity across both low- and high-risk strata. To further evaluate the significance of these results across random seeds, we combined the p-values obtained from each seed using Fisher’s method [9]. The test statistic was computed as $X^{2}=-2\sum_{i=1}^{k}\ln(p_{i})$, which follows a $\chi^{2}$ distribution with $2k$ degrees of freedom. Detailed p-values for each comparison are provided in Supplementary Table B.1 and Supplementary Table B.2.
Figure 4: Model survival analysis performance under the IECV setting. Regression analysis in the interval-external cross-validation setting, with model-predicted inverse risk (normalized) on the horizontal axis versus ground-truth survival time (normalized) on the vertical axis. Top panels represent internal validation and bottom panels show external validation in correlated cohorts. Cases with MACE=0 are denoted in black, while those with MACE=1 are indicated in blue.
As illustrated in Fig. 4, the linear relationship between log-transformed inverse risk scores from our prognostic model and observed survival times in right-censored specimens further confirms the discriminative efficacy of computed risk in characterizing survival distributions. Specimens with MACE=1 consistently occupy high-risk positions across test sets in IECV setting. Critically, validation under external-center training conditions manifested only a minor slope reduction in average relative to internal validation. Specifically, the confidence band in the external GLCCM cohort exhibits progressive convergence with extended survival times, yielding statistically narrower 95% confidence intervals versus the green internal band. This demonstrates time-invariant precision in survival estimation and temporal stability within the IECV experimental setting.